Weighted Average???

을n ucation.com apter 5 Problems Seved 5 Montoure Company uses a perpetual inventory system. It entered into the following calendar-year purchases and sales trans Units Sold at Retail DateActivities Jan. 1 Beginning inventory Feb. 10 Purchase Mar. 13 Purchase Mar. 15 Sales Aug. 21 Purchase Sept. 5 Purchase Sept. 10 Sales Units Acquired at Cost 600 units $60 per unit 480 units $57 per unit 120 units $42 per unit 5 nts 785 units $80 per unit 180 units $65 per unit 470 units $63 per unit 01:39:05 650 units $80 per unit Totals 1,850 units 1,435 units eBook Required: 1. Compute cost of goods available for sale and the number of units available for sale. Print Cost of goods available for sale $109.71 Number of units available for sale1.850 units eferences 2. Compute the number of units in ending inventory.
media%2F52b%2F52b80d5d-81d0-4e0b-a07e-bc
Answer #1

Answer:

1

cost of goods avialable
for sale

109710

Units   avalable for sales

1850

Working notes for the above answer:

Particular

Units
avalable for sales

Cost
per Unit

cost of goods avialable
for sale

A

B

C=A*B

beg.

600

60

36000

Add:Purchase

Feb.10

480

57

27360

Mar.13

120

42

5040

Aug.21

180

65

11700

Sep.5

470

63

29610

1850

109710

___________________________________

2

Weighted average perpetual

Goods purchasesd

Cost of goods sold

Inventory
balance

Date

Number
of units

at

Cost
per unit

Number
of units

at

Cost
per unit

Cost of
goods sold

Number
of units

at

Cost
per unit

Inventory
balance

Jan.1

600

60

600

60

36000

Feb.10

480

57

600

60

36000

480

57

27360

1080

58.67

63360

Weighted average cost =63360/1080 =58.67

Mar.13

120

42

600

60

36000

480

57

27360

120

42

5040

1200

57.00

68400

Weighted average cost =68400/1200 =57

Mar.15

785

57.00

44745

415

57.00

23655

Aug.21

180

65

415

57

23655

180

65

11700

595

59.42

35355

Weighted average cost =36355/595=59.42

415

57

23655

Sep.5

470

63

180

65

11700

470

63

29610

1065

61.00

64965

Weighted average cost =64965/1065=61

Sep.10

650

61.00

39650

415

25315

Total

84395

25315